# Tallahassee, FL 32306-4350, USA

### v1

@@ 3 Constraints on new physics from current data
We perform a global fit of the SMEFT including EW, Drell-Yan, Higgs-boson, top-quark, and flavour observables. The fit has been implemented in the NPSMEFTd6General model of the open-source HEPfit framework DeBlas:2019ehy ; hepfitsite . In this section, after specifying the adopted statistical framework and fitting procedure, we describe the methods used for the calculation of the different classes of observables entering the fit and provide references for the corresponding experimental measurements.
@@ 3.1 Introduction to the fitting procedure
We perform a Bayesian analysis to extract the posterior probability density for the SMEFT coefficients $C_{i}(\Lambda)$ given the experimental data, a uniform prior for $C_{i}(\Lambda)$ in the $[-4\pi,4\pi]\sim[-15,15]$ perturbative range [^4] This limit can be understood by demanding that a loop contribution from dimension-6 operators gives a contribution not larger than a genuine dimension-eight operator, and so forth for higher and higher dimensions. This is satisfied for $C_{i}\lesssim 4\pi$, following naive dimensional analysis Manohar:1983md ; Manohar:2018aog ., and a Gaussian prior for the input parameters $G_{F}$, $\alpha$ or $M_{W}$ (depending on the input scheme), $M_{Z}$, $M_{h}$, $\alpha_{s}(M_{Z})$ and the lepton masses, as described in Section 2.2. Concerning instead quark masses and CKM matrix elements, we consider uniform priors in the ranges reported in Table 2, and use the experimental information to constrain the relevant combination of SM parameters and SMEFT cofficients. SM parameters are always floating in the fit, while for SMEFT coefficients we perform fits with one operator at a time and fits with all coefficients floating simultaneously. To assess the impact of the different sets of observables, we also perform fits excluding one set at a time, and determine in this way which observables gives the strongest constraint. For each fit, the probability density function (pdf) of the SMEFT coefficients $C_{i}(\Lambda)$ is sampled using its numerical representation through the Markov Chain Monte Carlo (MCMC) implemented in HEPfit using the BAT library Caldwell:2008fw . From the MCMC samples, we compute $68\%$ and $95\%$ Highest Posterior Density Intervals (HPDI) for the SMEFT coefficients. From the largest absolute value of the $95\%$ HPDI we compute the lower bound on the effective NP interaction scale $\Lambda/\sqrt{|C|}$.
@@
Table 2: Ranges of the SM-parameter uniform priors used in the fits. The values of the quark masses are given at the corresponding renormalization scale, as indicated in the table. The CKM parameters are defined in the Wolfenstein parametrisation Wolfenstein:1983yz .
@@ 3.2 General setup for theoretical predictions
where SMEFT linear contributions come from the linear-order terms in both the $\delta{\cal O}_{\mathrm{SMEFT}}(\{p\})$ direct contribution and from expressing $p=p(\{\tilde{p}\})$ in terms of measured input parameters in ${\cal O}_{\mathrm{SM}}(\{p\})$. Choosing a given input scheme modifies the dependence of ${\cal O}(\{\tilde{p}\})$ on SMEFT coefficients. We note that for the purpose of illustration all the results presented in Section 4 have been obtained in the $M_{W}$-scheme, which is commonly used for global fits that include a majority of collider observables. 666 The only exceptions are the predictions for LHC Drell-Yan observables which, as will be explained below, are computed with the HighPT package Allwicher:2022mcg . These are given in the $\alpha$-scheme. Given that the difference between the $\alpha$ and $M_{W}$ schemes involves only operators that are well constrained by EW and Higgs precision observables, while the high-energy Drell-Yan measurements target four-fermion interactions, we expect this scheme mismatch to have a very small effect in the results.
Using Eq. (4), for all the observables considered in our study we have implemented state-of-the-art SM predictions (including QCD and EW corrections that could be analytically or numerically parametrised in an efficient way in the code) while SMEFT corrections, obtained from the interference between SMEFT and SM amplitudes, have been calculated at the lowest perturbative order. For several observables, e.g. the high-pT processes in top-quark physics, these are rescaled by the effect of higher-order SM corrections, assuming that their effects in the SM and the SMEFT are similar. When needed, further details will be given in the following sections where we discuss specific sets of observables.
@@ 3.3 Electroweak and Drell-Yan observables
We include in the fits the whole set of EW precision observables measured at the $Z$ pole at LEP/SLC. [^7] We use the results obtained without assuming lepton universality. A thorough description of all elements entering the EW global fit, including references to the state-of-the-art theoretical parametrisations used in HEPfit, is given in refs. deBlas:2021wap ; deBlas:2022hdk , to which we refer the reader interested in such details.
[…]
@@
Table 3: Observables relevant for the determination of the SM input parameters used in the fits (in the $W$ mass scheme). The values of the quark masses are taken from FLAG averages FLAG:2021npn . For $\alpha_{s}(M_{Z})$ we use the recent lattice QCD result of ref. Brida:2025gii . See the text for details on the $W$ mass and width.
Measurements of the $W$ branching ratios are also included, both from LEP2 ALEPH:2013dgf and from the recent (and more precise) determination by CMS CMS:2022mhs . We also include the measurements available for ratios of $W$ and $Z$ decays into different lepton flavours from Tevatron D0:1999rsi ; CDF:2005bdv ; D0:1999bqi ; CDF:1991mse and LHC LHCb:2016zpq ; ATLAS:2016nqi ; ATLAS:2020xea ; LHCb:2018ogb , which serve as tests of lepton flavour universality.
Two-to-two fermion processes above the $Z$ pole are also considered in our analysis. First, we include the measurements of total cross section and asymmetries for the $e^{+}e^{-}\to\mu^{+}\mu^{-},\tau^{+}\tau^{-}$ channels, the hadronic cross section, and the differential cross sections in $e^{+}e^{-}\to e^{+}e^{-}$ taken from LEP2 ALEPH:2013dgf  [^10] On the hadronic side, we do not include the measurements of $R_{b}$ and $R_{c}$, since each is derived assuming the other is SM-like, which in general will not be true in the SMEFT. Di-lepton ($\ell^{+}\ell^{-}$) and single-lepton ($\ell\nu$) processes from LHC Drell-Yan measurements are also included, implementing in HEPfit the predictions from HighPT Allwicher:2022mcg . These two-to-two fermion processes measured at high energies are largely sensitive to four-fermion operators, since their effects grow with the energy.
Finally, we also include in our study the constraints from $W^{+}W^{-}$ production at LEP2. For this observable we take into account the information available for the inclusive and differential cross sections measured at energies between 189 and 208 GeV, following the study in Berthier:2016tkq .
@@ 3.4 Higgs-boson observables
We consider the full combination of Higgs-boson signal strengths at 8 TeV from ATLAS and CMS as presented in ref. ATLAS:2016neq as well as the most recent combined measurements of Higgs-boson production cross sections, decay branching ratios, and simplified template cross sections (STXS) as provided by the ATLAS and CMS collaborations using 139 fb-1 of collision data at $\sqrt{s}=13$ TeV ATLAS-CONF-2021-053 ; ATLAS:2022vkf ; ATLAS:2024lyh ; ATLAS:2024fkg ; CMS:2022dwd ; CMS-PAS-HIG-21-018 . All measurements are separated by production and decay modes. These include measurements of Higgs-boson production via gluon-gluon fusion (ggF), vector-boson fusion (VBF), and associated production with gauge bosons ($ZH$ and $WH$) and top quarks ($t\bar{t}H$ and $tH$) as well as measurements of Higgs-boson decays into $\gamma\gamma,ZZ^{*},WW^{*},\tau^{+}\tau^{-},b\bar{b},\mu^{+}\mu^{-}$, and $Z\gamma$. For the STXS we adopt the most recent Stage 1.2 binning and compare with the experimental results presented in ATLAS:2024lyh ; CMS-PAS-HIG-21-018 . In all cases we include the full correlation matrices provided by the experimental collaborations.
@@
We parametrise both cross sections and decay rates at linear order in the SMEFT coefficients using MG5_aMC@NLO Alwall:2014hca with a UFO file developed for this study. [^11] The UFO file used for this study has been generated with the SMEFTci2 Mathematica in-house code. This is based on SMEFTsim Brivio:2017btx ; Brivio:2020onw and has also been validated against SmeftFR v3 Dedes:2023zws . To obtain the parametrisation of some SM loop-induced processes, e.g. ggF, we also use the SMEFT@NLO package Degrande:2020evl .
@@
As mentioned earlier, we work in the $M_{W}$-scheme and use as default input parameters the experimental values of $G_{F}$, $M_{Z}$, and $M_{W}$. We adopt the NNPDF4.0 NLO set of parton distribution functions with $\alpha_{s}(M_{Z})=0.118$ and fix factorisation and renormalisation scales to the relevant scale for each process (e.g. $\mu=M_{H}$ for ggF, etc.). All chosen scales are very close to $\mu_{W}$, the scale at which we evaluate the RG-evolved SMEFT Wilson coefficients such that, for the purpose of this study, the effects of running from $\mu_{W}$ to the actual scale at which the Higgs-boson production processes have been calculated is negligible. Likewise, the theoretical error induced by the choice of PDF and their variation (at the few per-cent level for most state-of-the-art SM calculations) is also below the sensitivity of our study.
@@
The parametrisation of the STXS has been obtained at parton level. For the purpose of the SMEFT global fit, working at parton level captures the bulk of the relevant SMEFT effects. Indeed, we did not notice any substantial difference from the parametrisations reported e.g. in ATLAS-CONF-2021-053 where parton-shower effects have been fully simulated, except for a few bins that could not be obtained at parton level. This includes for instance the case of ggF Higgs-boson production with 0 jets (where the STXS bin is defined by a cut on the Higgs-boson transverse momentum) or the case of Higgs-boson plus two jets final states (where the STXS bins are defined by a cut on the transverse momenta of the whole final-state system) since in both cases the transverse momentum used to define the STXS bin cannot be properly defined without considering extra final-state partons as produced by the parton-shower. In these cases we have removed the corresponding bin.
@@ 3.5 Top-quark observables
Process Observable $\sqrt{s}$$\cal{L}$ SM Ref.
$p\bar{p}\rightarrow t\bar{t}$ $\rm{d}A^{t\bar{t}}_{\rm{FB}}/\rm{d}m_{t\bar{t}}$ 1.96 TeV 9.7 fb-1 Czakon:2014xsa ; Czakon:2016ckf CDF:2017cvy
$pp\rightarrow t\bar{t}$ $\sigma_{t\bar{t}}^{13\rm{TeV}}/\sigma_{t\bar{t}}^{8\rm{TeV}}$ 13 & 8 TeV 20 & 36 fb-1 Czakon:2011xx ; Czakon:2013goa ATLAS:2019hau
$\sigma_{t\bar{t}}^{8\rm{TeV}}/\sigma_{t\bar{t}}^{7\rm{TeV}}$ 8 & 7 TeV 20 & 5 fb-1 ATLAS:2022aof
$\sigma_{t\bar{t}}$ 13 TeV 36/139 fb-1 ATLAS:2019hau ; ATLAS:2020aln
$\rm{d}\sigma_{t\bar{t}}/\rm{d}m_{t\bar{t}}$ 13 TeV 36 fb-1 ATLAS:2019hxz
$(\rm{d}\sigma_{t\bar{t}}/\rm{d}m_{t\bar{t}})/\sigma_{t\bar{t}}$ 13 TeV 36/137 fb-1 CMS:2018adi ; CMS:2021vhb
$\rm{d}A_{C}/\rm{d}m_{t\bar{t}}$ 13 TeV 140 fb-1 ATLAS:2022waa
$pp\rightarrow t\bar{t}Z$ $d\sigma/dp_{T}^{Z}$ 13 TeV 77.5/140 fb-1 Broggio:2019ewu CMS:2019too /ATLAS:2023eld
$pp\rightarrow t\bar{t}\gamma$ $d\sigma/dp_{T}^{\gamma}$ 13 TeV 140 fb-1 Bevilacqua:2018woc ; Bevilacqua:2018dny CMS:2021klw ; CMS:2022lmh ; ATLAS:2020yrp
$pp\rightarrow t\bar{t}W$ $\sigma_{ttW^{\pm}}$ 13 TeV 140 fb-1 Frederix:2017wme ; Buonocore:2023ljm ATLAS:2024moy ; CMS:2022tkv
@@
$t\rightarrow Wb$ $F_{0}$, $F_{L}$ 8 TeV 20 fb-1 PhysRevD.81.111503 Aad:2020jvx
13 TeV 140 fb-1 ATLAS:2022rms
$pp\rightarrow tW$ $\sigma$ 7 TeV 4.6 & 1.5 fb-1 Kidonakis:2010ux ATLAS:2019hhu
8 TeV 20 fb-1 Kidonakis:2010ux ATLAS:2019hhu
13 TeV 3.2/140 fb-1 Kidonakis:2010ux ATLAS:2016ofl /CMS:2022ytw
$pp\rightarrow t\bar{b}$ (s-ch) $\sigma$ 8 TeV 20 fb-1 Aliev:2010zk ; Kant:2014oha ATLAS:2019hhu
13 TeV 140 fb-1 Muller:2723842 ; ATLAS:2022wfk
$pp\rightarrow tq$ (t-ch) $\sigma$ 7 TeV 4.6 & 1.5 fb-1 Aliev:2010zk ; Kant:2014oha ATLAS:2019hhu
8 TeV 20 fb-1 ATLAS:2019hhu
13 TeV 36/140 fb-1 CMS:2018lgn /ATLAS:2024ojr
$pp\rightarrow t\gamma q$ $\sigma$ 13 TeV 140/36 fb-1 Alwall:2014hca ATLAS:2023qdu /CMS:2018hfk
$pp\rightarrow tZq$ $\sigma$ 13 TeV 140 fb-1 Alwall:2014hca CMS:2021ugv ; ATLAS:2020bhu
$pp\rightarrow t\bar{t}b\bar{b}$ $\sigma$ 13 TeV 36 fb-1 Alwall:2014hca CMS:2020grm
$pp\rightarrow t\bar{t}t\bar{t}$ $\sigma$ 13 TeV 140 fb-1 vanBeekveld:2022hty ATLAS:2021kqb ; CMS:2023ftu ; ATLAS:2023ajo

### v2

@@ 3 Constraints on new physics from current data
We perform a global fit of the SMEFT including EW, Drell-Yan, Higgs-boson, top-quark, and flavour observables. The fit has been implemented in the NPSMEFTd6General model of the open-source HEPfit framework J. de Blas et al. (2020); 1. In this section, after specifying the adopted statistical framework and fitting procedure, we describe the methods used for the calculation of the different classes of observables entering the fit and provide references for the corresponding experimental measurements.
@@ 3.1 Introduction to the fitting procedure
We perform a Bayesian analysis to extract the posterior probability density for the SMEFT coefficients $C_{i}(\Lambda)$ given the experimental data, a uniform prior for $C_{i}(\Lambda)$ in the $4\pi,4\pi]\sim[-15,15]$ perturbative range [^4] This limit can be understood by demanding that a loop contribution from dimension-6 operators gives a contribution not larger than a genuine dimension-eight operator, and so forth for higher and higher dimensions. This is satisfied for $C_{i}\lesssim 4\pi$, following naive dimensional analysis Manohar and Georgi (1984); Manohar (2018)., and a Gaussian prior for the input parameters $G_{F}$, $\alpha$ or $M_{W}$ (depending on the input scheme), $M_{Z}$, $M_{h}$, $\alpha_{s}(M_{Z})$ and the lepton masses, as described in Section 2.2. Concerning instead quark masses and CKM matrix elements, we consider uniform priors in the ranges reported in Table 2, and use the experimental information to constrain the relevant combination of SM parameters and SMEFT cofficients. SM parameters are always floating in the fit, while for SMEFT coefficients we perform fits with one operator at a time and fits with all coefficients floating simultaneously. To assess the impact of the different sets of observables, we also perform fits excluding one set at a time, and determine in this way which observables gives the strongest constraint. For each fit, the probability density function (pdf) of the SMEFT coefficients $C_{i}(\Lambda)$ is sampled using its numerical representation through the Markov Chain Monte Carlo (MCMC) implemented in HEPfit using the BAT library Caldwell et al. (2009). From the MCMC samples, we compute $68\%$ and $95\%$ Highest Posterior Density Intervals (HPDI) for the SMEFT coefficients. From the largest absolute value of the $95\%$ HPDI we compute the lower bound on the effective NP interaction scale $\Lambda/\sqrt{|C|}$.
@@
Table 2: Ranges of the SM-parameter uniform priors used in the fits. The values of the quark masses are given at the corresponding renormalization scale, as indicated in the table. The CKM parameters are defined in the Wolfenstein parametrisation Wolfenstein (1983).
@@ 3.2 General setup for theoretical predictions
[…]
Using Eq. (4), for all the observables considered in our study we have implemented state-of-the-art SM predictions (including QCD and EW corrections that could be analytically or numerically parametrised in an efficient way in the code) while SMEFT corrections, obtained from the interference between SMEFT and SM amplitudes, have been calculated at the lowest perturbative order. For several observables, e.g. the high-$p_{T}$ processes in top-quark physics, these are rescaled by the effect of higher-order SM corrections, assuming that their effects in the SM and the SMEFT are similar. When needed, further details will be given in the following sections where we discuss specific sets of observables.
@@ 3.3 Electroweak and Drell-Yan observables
We include in the fits the whole set of EW precision observables measured at the $Z$ pole at LEP/SLC. [^7] We use the results obtained without assuming lepton universality. A thorough description of all elements entering the EW global fit, including references to the state-of-the-art theoretical parametrisations used in HEPfit, is given in refs. de Blas et al. (2022a, b), to which we refer the reader interested in such details.
For the $W$ mass we consider the inputs from LEP2, ATLAS, CMS, LHCb and Tevatron, excluding the recent CDF measurement Aaltonen and others (2022) which, in the light of the very recent CMS result Chekhovsky and others (2024), requires additional investigation. [^8] In ref. de Blas et al. (2022b) we discussed the impact of the CDF measurement on the EW precision fit, and introduced a conservative scenario in which the uncertainty on the $M_{W}$ average including CDF was inflated to ensure compatibility among experimental measurements. This is less justified after the very recent CMS result, whose accuracy is comparable to the CDF one but is fully consistent with all existing measurements except for the CDF one. We combine all the hadronic $W$-mass measurements obtained using the CT18 PDF set, assuming the same PDF correlation indicated in Amoroso and others (2024), which we take to be the same for ATLAS and CMS (not present in that reference). We also assume that QED systematics are 100% correlated across the different measurements. The hadron-collider result is then combined with the LEP2 $W$-mass determination in an uncorrelated manner. This results in an average value $M_{W}=80.3635\pm 0.0080$ GeV. Similarly, we combine the measurements of the $W$ width results from LEP2, ATLAS and Tevatron. There is some tension between the ATLAS and Tevatron determinations ($\chi^{2}/\textrm{d.o.f.}\approx 3$), which we take into account by scaling the uncertainty of the combination, before adding the LEP2 result. Finally, since the measurements of the $W$ width are not available for all experiments providing results for $M_{W}$, we neglect the correlation between the two observables. We obtain $\Gamma_{W}=2.151\pm 0.049$ GeV. For the $Z$-boson mass, we also take into account the CDF determinations in ref. Aaltonen and others (2022), in good agreement with the LEP value. The Higgs-boson mass completes the list of SM inputs needed in the $W$-mass scheme, and here we use the same value as in ref. de Blas et al. (2022b). The numerical values used in the fit for those observables directly related to the SM inputs are reported in Table 3.999The Fermi constant $G_{F}$, known to a significantly higher precision, is fixed in the fits.
@@
Table 3: Observables relevant for the determination of the SM input parameters used in the fits (in the $W$ mass scheme). The values of the quark masses are taken from FLAG averages Aoki and others (2022). For $\alpha_{s}(M_{Z})$ we use the recent lattice QCD result of ref. Brida et al. (2025). See the text for details on the $W$ mass and width.
Measurements of the $W$ branching ratios are also included, both from LEP2 Schael and others (2013) and from the recent (and more precise) determination by CMS Tumasyan and others (2022d). We also include the measurements available for ratios of $W$ and $Z$ decays into different lepton flavours from Tevatron Abbott and others (2000b); Abulencia and others (2007); Abbott and others (2000a); Abe and others (1992) and LHC Aaij and others (2016); Aaboud and others (2017); Aad and others (2021b); Aaij and others (2018), which serve as tests of lepton flavour universality.
Two-to-two fermion processes above the $Z$ pole are also considered in our analysis. First, we include the measurements of total cross section and asymmetries for the $e^{+}e^{-}\to\mu^{+}\mu^{-},\tau^{+}\tau^{-}$ channels, the hadronic cross section, and the differential cross sections in $e^{+}e^{-}\to e^{+}e^{-}$ taken from LEP2 Schael and others (2013) [^10] On the hadronic side, we do not include the measurements of $R_{b}$ and $R_{c}$, since each is derived assuming the other is SM-like, which in general will not be true in the SMEFT. Di-lepton ($\ell^{+}\ell^{-}$) and single-lepton ($\ell\nu$) processes from LHC Drell-Yan measurements are also included, implementing in HEPfit the predictions from HighPT Allwicher et al. (2023b). These two-to-two fermion processes measured at high energies are largely sensitive to four-fermion operators, since their effects grow with the energy.
Finally, we also include in our study the constraints from $W^{+}W^{-}$ production at LEP2. For this observable we take into account the information available for the inclusive and differential cross sections measured at energies between 189 and 208 GeV, following the study in Berthier et al. (2016).
@@ 3.4 Higgs-boson observables
We consider the full combination of Higgs-boson signal strengths at 8 TeV from ATLAS and CMS as presented in ref. Aad and others (2016) as well as the most recent combined measurements of Higgs-boson production cross sections, decay branching ratios, and simplified template cross sections (STXS) as provided by the ATLAS and CMS collaborations using 139 fb-1 of collision data at $\sqrt{s}=13$ TeV 77; G. Aad et al. (2022); G. Aad et al. (2024c); G. Aad et al. (2025); A. Tumasyan et al. (2022a); 76. All measurements are separated by production and decay modes. These include measurements of Higgs-boson production via gluon-gluon fusion (ggF), vector-boson fusion (VBF), and associated production with gauge bosons ($ZH$ and $WH$) and top quarks ($t\bar{t}H$ and $tH$) as well as measurements of Higgs-boson decays into $\gamma\gamma,ZZ^{*},WW^{*},\tau^{+}\tau^{-},b\bar{b},\mu^{+}\mu^{-}$, and $Z\gamma$. For the STXS we adopt the most recent Stage 1.2 binning and compare with the experimental results presented in G. Aad et al. (2024c); 76. In all cases we include the full correlation matrices provided by the experimental collaborations.
@@
We parametrise both cross sections and decay rates at linear order in the SMEFT coefficients using MG5_aMC@NLO Alwall et al. (2014) with a UFO file developed for this study. [^11] The UFO file used for this study has been generated with the SMEFTci2 Mathematica in-house code. This is based on SMEFTsim Brivio et al. (2017); Brivio (2021) and has also been validated against SmeftFR v3 Dedes et al. (2024). To obtain the parametrisation of some SM loop-induced processes, e.g. ggF, we also use the SMEFT@NLO package Degrande et al. (2021).
@@
As mentioned earlier, we work in the $M_{W}$-scheme and use as default input parameters the experimental values of $G_{F}$, $M_{Z}$, and $M_{W}$. We adopt the NNPDF4.0 NLO set of parton distribution functions with $\alpha_{s}(M_{Z})=0.118$ and fix factorisation and renormalization scales to the relevant scale for each process (e.g. $\mu=M_{H}$ for ggF, etc.). All chosen scales are very close to $\mu_{W}$, the scale at which we evaluate the RG-evolved SMEFT Wilson coefficients such that, for the purpose of this study, the effects of running from $\mu_{W}$ to the actual scale at which the Higgs-boson production processes have been calculated is negligible, specially given the experimental uncertainty on the high-$p_{T}$ bins. [^12] As shown in Ref. Battaglia et al. (2021) (also recently revisited in Ref. Maltoni et al. (2024)) these effects are not critical given the current experimental precision but may be relevant in the near future when data from HL-LHC becomes available. Likewise, the theoretical error induced by the choice of PDF and their variation (at the few per-cent level for most state-of-the-art SM calculations) is also below the sensitivity of our study.
@@
The parametrisation of the STXS has been obtained at parton level. For the purpose of the SMEFT global fit, working at parton level captures the bulk of the relevant SMEFT effects. Indeed, we did not notice any substantial difference from the parametrisations reported e.g. in 77 where parton-shower effects have been fully simulated, except for a few bins that could not be obtained at parton level. This includes for instance the case of ggF Higgs-boson production with 0 jets (where the STXS bin is defined by a cut on the Higgs-boson transverse momentum) or the case of Higgs-boson plus two jets final states (where the STXS bins are defined by a cut on the transverse momenta of the whole final-state system) since in both cases the transverse momentum used to define the STXS bin cannot be properly defined without considering extra final-state partons as produced by the parton-shower. In these cases we have removed the corresponding bin.
@@ 3.5 Top-quark observables
Process Observable $\sqrt{s}$ $\int\cal{L}$ SM Ref.
$p\bar{p}\rightarrow t\bar{t}$ $\rm{d}A^{t\bar{t}}_{\rm{FB}}/\rm{d}m_{t\bar{t}}$ 1.96 TeV 9.7 fb-1 Czakon et al. (2015, 2016) Aaltonen and others (2018)
$pp\rightarrow t\bar{t}$ $\sigma_{t\bar{t}}^{13\rm{TeV}}/\sigma_{t\bar{t}}^{8\rm{TeV}}$ 13 & 8 TeV 20 & 36 fb-1 Czakon and Mitov (2014); Czakon et al. (2013) Aad and others (2020a)
$\sigma_{t\bar{t}}^{8\rm{TeV}}/\sigma_{t\bar{t}}^{7\rm{TeV}}$ 8 & 7 TeV 20 & 5 fb-1 Aad and others (2023a)
$\sigma_{t\bar{t}}$ 13 TeV 36/139 fb-1 Aad and others (2020a, b)
$\rm{d}\sigma_{t\bar{t}}/\rm{d}m_{t\bar{t}}$ 13 TeV 36 fb-1 Aad and others (2019)
$(\rm{d}\sigma_{t\bar{t}}/\rm{d}m_{t\bar{t}})/\sigma_{t\bar{t}}$ 13 TeV 36/137 fb-1 Sirunyan and others (2019); Tumasyan and others (2021a)
$\rm{d}A_{C}/\rm{d}m_{t\bar{t}}$ 13 TeV 140 fb-1 Aad and others (2023b)
$pp\rightarrow t\bar{t}Z$ $d\sigma/dp_{T}^{Z}$ 13 TeV 77.5/140 fb-1 Broggio et al. (2019) Sirunyan and others (2020c)/Aad and others (2024b)
$pp\rightarrow t\bar{t}\gamma$ $d\sigma/dp_{T}^{\gamma}$ 13 TeV 140 fb-1 Bevilacqua et al. (2018, 2019) Tumasyan and others (2021b, 2022c); Aad and others (2020c)
$pp\rightarrow t\bar{t}W$ $\sigma_{ttW^{\pm}}$ 13 TeV 140 fb-1 Frederix et al. (2018); Buonocore et al. (2023) Aad and others (2024e); Tumasyan and others (2023b)
@@
$t\rightarrow Wb$ $F_{0}$, $F_{L}$ 8 TeV 20 fb-1 Czarnecki et al. (2010) ATLAS and CMS Collaborations (2020)
13 TeV 140 fb-1 Aad and others (2023d)
$pp\rightarrow tW$ $\sigma$ 7 TeV 4.6 & 1.5 fb-1 Kidonakis (2010) Aaboud and others (2019)
8 TeV 20 fb-1 Kidonakis (2010) Aaboud and others (2019)
13 TeV 3.2/140 fb-1 Kidonakis (2010) Aaboud and others (2018)/Tumasyan and others (2023a)
$pp\rightarrow t\bar{b}$ (s-ch) $\sigma$ 8 TeV 20 fb-1 Aliev et al. (2011); Kant et al. (2015) Aaboud and others (2019)
13 TeV 140 fb-1 Muller (2020); Aad and others (2023c)
$pp\rightarrow tq$ (t-ch) $\sigma$ 7 TeV 4.6 & 1.5 fb-1 Aliev et al. (2011); Kant et al. (2015) Aaboud and others (2019)
8 TeV 20 fb-1 Aaboud and others (2019)
13 TeV 36/140 fb-1 Sirunyan and others (2020b)/Aad and others (2024d)
$pp\rightarrow t\gamma q$ $\sigma$ 13 TeV 140/36 fb-1 Alwall et al. (2014) Aad and others (2023f)/Sirunyan and others (2018)
$pp\rightarrow tZq$ $\sigma$ 13 TeV 140 fb-1 Alwall et al. (2014) Tumasyan and others (2022b); Aad and others (2020d)
$pp\rightarrow t\bar{t}b\bar{b}$ $\sigma$ 13 TeV 36 fb-1 Alwall et al. (2014) Sirunyan and others (2020a)
$pp\rightarrow t\bar{t}t\bar{t}$ $\sigma$ 13 TeV 140 fb-1 van Beekveld et al. (2023) Aad and others (2021a); Hayrapetyan and others (2023); Aad and others (2023e)